OpenStudy (symone19):

Convert 72°F to °C (to the nearest tenth) using one of the following formulas: F = 9/5xC + 32 or C = (F — 32) x 5/9. 22.2 54.2 72.0 161.6

OpenStudy (shadowlegendx):

Which one do you want to use?

OpenStudy (shadowlegendx):

\[C = (F - 32) \times \frac{ 5 }{ 9 }\] It clearly shows what we have to do. Lets start by inputting F = 72 \[C = (72 - 32) \times \frac{ 5 }{ 9 }\] Do the functions inside the parentheses(subtraction) \[C = 40 \times \frac{ 5 }{ 9 }\] Multiply 40 by 5/9 We get, \[C = 22.222222222 \rightarrow C = 22.2\]